### v1

# Succeeding at Scale: Automated Multi-Retriever Fusion and Query-Side Adaptation for Multi-Tenant Search

Abstract
Large-scale multi-tenant retrieval systems amass vast user query logs yet critically lack the curated relevance labels required for effective domain adaptation. This "dark data" problem is exacerbated by the operational cost of model updates: jointly fine-tuning query and document encoders requires re-indexing the entire corpus, which is prohibitive in multi-tenant environments with thousands of isolated indices. To address these dual challenges, we introduce DevRev Search, a passage retrieval benchmark for technical customer support constructed through a fully automatic pipeline. We employ a fusion-based candidate generation strategy, pooling results from diverse sparse and dense retrievers, and utilize an LLM-as-a-Judge to perform rigorous consistency filtering and relevance assignment. We further propose a practical Index-Preserving Adaptation strategy: by fine-tuning only the query encoder via Low-Rank Adaptation (LoRA), we achieve competitive performance improvements while keeping the document index frozen. Our experiments on DevRev Search and SciFact demonstrate that targeting specific transformer layers in the query encoder yields optimal quality-efficiency trade-offs, offering a scalable path for personalized enterprise search.
Succeeding at Scale: Automated Multi-Retriever Fusion and Query-Side Adaptation for Multi-Tenant Search
Prateek Jain1thanks: Corresponding author: prateek.jain@devrev.ai, Shabari S Nair2, Ritesh Goru1, Prakhar Agarwal3,
Ajay Yadav2, Yoga Sri Varshan Varadharajan2, Constantine Caramanis2 1DevRev, Austin, USA 3DevRev, Bengaluru, India 2The University of Texas at Austin, Austin, USA
1 Introduction
The transition from lexical matching (e.g., BM25 Robertson and Zaragoza (2009)) to dense neural retrieval has revolutionized information discovery Karpukhin et al. (2020). However, deploying bi-encoder architectures in multi-tenant enterprise environments presents a "double scarcity" challenge. First, the Data Scarcity Bottleneck: enterprise tenants possess "dark data," proprietary corpora where relevance labels are non-existent, and standard benchmarks like BEIR Thakur et al. (2021) fail to capture the noisy, heterogeneous nature of these domains. Second, the Adaptation Latency Bottleneck: symmetric fine-tuning of both encoders Qu et al. (2021) incurs a massive "Re-indexing Tax," as any document encoder update necessitates re-generating embeddings for the entire corpus, computationally prohibitive for platforms hosting thousands of tenants.
To bridge these gaps, we present a unified methodology for scalable dataset construction and efficient model adaptation:
  1. 1.
    DevRev Search Benchmark: A scalable pipeline that synthesizes training data without human annotators by pooling candidates from diverse retrievers via Reciprocal Rank Fusion Cormack et al. (2009) and employing LLM-as-a-Judge filtering Dai et al. (2023); Rahmani et al. (2024).
  2. 2.
    Zero-Reindexing Adaptation: An asymmetric fine-tuning strategy adapting only the query encoder via LoRA Hu et al. (2021), enabling tenant-specific adapters on a shared, frozen document index.
  3. 3.
    Layer Sensitivity Analysis: Empirical evidence that targeting specific transformer layers in the query encoder maximizes Recall while minimizing trainable parameters.
We validate our approach on DevRev Search and SciFact Wadden et al. (2022), demonstrating robust domain adaptation with a fraction of the cost of full fine-tuning.
2 Related Work
Synthetic Data.
Addressing label scarcity, recent works leverage LLMs for synthetic data generation Dai et al. (2023); Bonifacio et al. (2022); Wang et al. (2022). A key component is consistency filtering, discarding queries that fail to retrieve their source. We extend this by employing fusion-based candidate generation Cormack et al. (2009), aggregating diverse retrievers to minimize single-model bias and improve coverage.
LLM-as-Judge.
While LLMs show promise as relevance assessors Rahmani et al. (2024), concerns regarding bias persist Soboroff (2025). We mitigate this by using LLMs primarily for filtering pooled candidates, identifying positives rather than generating them, and validating a subset with human annotators.
Index-Preserving Adaptation.
Standard symmetric fine-tuning Karpukhin et al. (2020) incurs a prohibitive "Re-indexing Tax". Prior index-preserving approaches focus on pseudo-relevance feedback Yu et al. (2021) or asymmetric tuning Wang and Lyu (2023). We formalize this as Query-Side Adaptation, adapting the query manifold while freezing the document index.
Parameter-Efficient Fine-Tuning (PEFT).
PEFT methods like Adapters Houlsby et al. (2019) and LoRA Hu et al. (2021) have proven effective in retrieval Litschko et al. (2022). Unlike full fine-tuning, LoRA injects trainable low-rank matrices while freezing the backbone. Building on findings regarding intrinsic dimensionality Aghajanyan et al. (2020), we demonstrate that applying LoRA selectively to top query encoder layers maximizes efficiency for multi-tenant serving.
3 Dataset Generation
The recent paradigm shift in retrieval performance is largely attributed to the availability of high-quality, domain-specific data. However, a significant gap remains in the availability of publicly accessible enterprise search datasets that reflect the complex, semi-structured nature of real-world organizational data. Existing benchmarks like MS-MARCO or SciFact focus on web-scale passages or scientific abstracts, leaving a void for tasks involving technical support tickets, issue trackers, and internal documentation. By releasing the DevRev Search dataset, we aim to bridge this gap, providing the community with a high-fidelity benchmark for enterprise-specific retrieval.
Traditional manual annotation is not only prohibitively expensive and tedious but also suffers from low recall; human annotators cannot feasibly parse millions of documents, often leading to false negatives where relevant documents are overlooked simply because they were not reviewed. To address these challenges, we propose a scalable, automated pipeline to construct the DevRev Search dataset. Our methodology leverages a multi-stage process designed to maximize candidate coverage while maintaining high precision through the use of an LLM-as-judge.
3.1 Query Collection and Cleaning
We collected customer queries from production agent interactions as our source of real-world question data. However, raw customer queries often contain noise, including test queries, code snippets, and malformed inputs that are not legitimate natural language questions.
To ensure dataset quality, we implemented a multi-stage filtering process: (1) Length filtering: Removing queries with word counts in the bottom and top 25% percentiles. (2) Language detection: Retaining only English queries. (3) Deduplication: Removing exact duplicate queries. (4) Clustering-based diversity: Selecting representative samples from clusters to ensure diversity and avoid semantic repetition.
3.2 Multi-Retriever Annotation
To create high-quality query-document pairs, we employed an ensemble retrieval approach designed to maximize recall while maintaining precision.
Retrieval Ensemble: We applied an ensemble of seven diverse models: six dense retrievers (gemini-embedding-001 Google (2025), text-embedding-3-large OpenAI (2024), embed-english-v3 Cohere (2023), Qwen-3-Embedding-8B Zhang et al. (2025), GTE-Qwen2-7B-Instruct Li et al. (2023), SFR-Embedding-Mistral Meng et al. (2024)) and one lexical retriever (BM25), each returning the top 60 document chunks.
Union-based Aggregation: We computed the union of results from all retrievers to create a comprehensive candidate set of potentially relevant chunks. This union-based approach ensures that chunks retrieved by any of the seven models are included in the candidate set, maximizing coverage and recall across different retrieval paradigms yielding $\geq 60$ and $\leq 420$ unique candidate chunks per query across all models.
LLM-based Filtering: While the ensemble approach ensures high recall, it also introduces noise through false positives. To improve precision, we applied LLM-based filtering to the fused candidate set. Using a carefully designed prompt (provided in Appendix A.5), we tasked a large language model with identifying and retaining only the document chunks that genuinely contain information relevant to answering each query. This filtering step removes chunks that may have high lexical or semantic similarity to the query but lack substantive answer content.
Quality Validation: To verify the reliability of our automated annotation process, we randomly sampled $10\%$ queries and manually validated the final annotations. This validation confirmed the accuracy of our annotation pipeline.
See Appendix A for further details on document segmentation and dataset statistics.
4 Experiments and Results
Standard bi-encoder optimization updates both query ($E_{q}$) and document ($E_{d}$) encoders. However, modifying $E_{d}$ incurs a re-indexing tax - the need to re-embed millions of documents and rebuild high-dimensional vector indices (e.g., HNSW). In large-scale, multi-tenant systems, this compute-intensive process introduces significant downtime and synchronization latency. To maintain a high velocity of model improvement without the prohibitive cost of index reconstruction, we propose Query-Side Adaptation: freezing the document encoder and index to enable near-instantaneous deployment.
We evaluate this strategy on two contrasting domains. First, our DevRev Search dataset (enterprise) features high relevance density (avg. 13.6 relevant chunks/query), testing the model’s capacity for broad semantic coverage and high recall. Conversely, SciFact Wadden et al. (2022) (scientific) exhibits low density (avg. 1.1 relevant docs/query), requiring identification of unique, specific evidence with high precision. We employ snowflake-arctic-embed-l-v2.0 Yu et al. (2024) and Qwen3-Embedding-4B Zhang et al. (2025) backbones. Training optimizes InfoNCE loss with 8 mined hard negatives per query essential for distinguishing subtly different technical concepts using a cosine learning rate scheduler. Refer to A.6 for our detailed experimentation setup.
Our analysis covers: (1) Query-Only vs. Joint Tuning to quantify the performance trade-off of freezing the index; (2) LoRA Approximation and Scaling to analyze sensitivity to rank $r$; and (3) Module Targeting to identify which transformer components (e.g., Attention vs. MLP) yield the highest returns.
4.1 Comparison of Fine-Tuning Strategies
We first address whether query-encoder-only fine-tuning ($Q$) can achieve performance parity with joint Query-Document fine-tuning ($QD$). The results obtained by using the optimal hyperparameter settings for each configuration, are summarized in Fig 1.
DevRev Search Results.
On the DevRev Search benchmark surprisingly, Query-Only ($Q$) consistently outperforms Query-Document ($QD$) on the enterprise benchmark. We attribute this to asymmetric regularization. In low-resource specialized domains, joint optimization effectively doubles the parameter space, increasing the risk of overfitting and distorting the pre-trained document manifold. By freezing the document encoder, we enforce a structural constraint projecting queries into a stable target space which improves generalization to unseen test queries.
SciFact Results.
On the SciFact dataset, $QD$ remains the upper bound, confirming that complex scientific alignment benefits from reshaping both spaces. However, $Q$ remains highly competitive, recovering the vast majority of performance gains (within 1-2% of $QD$). This establishes Query-Only adaptation as a Pareto-optimal strategy for production: it delivers comparable accuracy to joint tuning while completely eliminating the prohibitive re-indexing tax.
(a) DevRev Search
(b) SciFact
Figure 1: Comparison of Recall@10 for Baseline, Query-Document (QDQD), and Query-Only ($Q$) fine-tuning. On DevRev Search, $Q$ surprisingly outperforms QDQD, while on SciFact, $Q$ remains highly competitive.
(a) DevRev Search
(b) SciFact
Figure 2: Performance vs. LoRA Rank. Note Qwen’s preference for lower ranks ($r=8$) on DevRev Search versus mid-ranks ($r=64$) on SciFact.
4.2 Impact of LoRA Rank
We next investigate the sensitivity of the models to the LoRA rank ($r$), which controls the capacity of the trainable adapters. As shown in Fig 2, the optimal rank for DevRev Search varies significantly by model architecture. In particular, since the DevRev Search dataset is relatively small, larger models like Qwen3-Embedding-4B tend to overfit on higher ranks. On SciFact however, we observe that higher ranks are generally preferred.
4.3 Targeted Lora Module Fine-tuning
Finally, we analyze which transformer sub-layers (Query-Value QV, Feed-Forward Network FFN, Query-Key-Value QKV, or All Layers) yield the best adaptation results.
(a) DevRev Search
(b) SciFact
Figure 3: Targeted Module Fine Tuning
The findings from Fig 3 highlight that the optimal fine-tuning strategy is highly dependent on the model-dataset pair. While the smaller Snowflake model benefits from maximizing capacity (Higher Rank, All Layers), the larger Qwen model performs well with targeted regularization (Lower Rank, QV/FFN modules) to achieve peak retrieval quality, indicating a clear trend.
5 Conclusion
We presented a unified approach to scalable dataset construction and efficient model adaptation for multi-tenant retrieval. Our automated pipeline, combining multi-retriever fusion with LLM-based consistency filtering, produces high-quality training data without manual annotation, yielding the DevRev Search benchmark. Our index-preserving adaptation strategy, fine-tuning only the query encoder via LoRA, achieves competitive performance while eliminating re-indexing overhead, with layer-targeted updates further optimizing the quality-efficiency trade-off.
Limitations
Our evaluation is limited to English queries across two domains; generalizability to other verticals and multilingual settings remains unexplored. Query-side adaptation may impose a performance ceiling compared to joint fine-tuning, and we do not investigate hybrid approaches with cross-encoder rerankers. Optimal LoRA configurations may vary across architectures, and our metrics do not directly measure downstream RAG performance.
Acknowledgments
We used AI writing assistants to assist with code development and manuscript refinement. All content was reviewed and verified by the authors.
Appendix
Appendix A Dataset Details
A.1 Document Segmentation and Semantic Granularity
Enterprise documentation is often characterized by high length and low information density relative to specific user intents. Large documents pose two primary challenges for dense retrieval: (1) Context Constraints, as most transformer-based encoders have a fixed token limit (e.g., 512 tokens), and (2) Information Compression, where the fixed-dimensional latent vector is insufficient to represent the entire semantic breadth of a long document, leading to "diluted" embeddings.
To mitigate these issues, we collected all public-facing documentation from DevRev and applied a Recursive Character Splitting strategy LangChain (2025). We chose a maximum chunk size of 500 characters with zero overlap to maximize the number of distinct semantic units. The recursive algorithm prioritizes splitting at logical boundaries (e.g., double newlines, then periods, then spaces) to maintain structural integrity. This ensures that each document fragment is small enough to be represented accurately by a single embedding while remaining large enough to contain a self-contained answer or concept.
A.2 Dataset Statistics
We partition the DevRev Search dataset into a training set and an evaluation (test) set. The dataset consists of 291 training queries and 92 test queries. The training set exhibits a rich density of relevant documents, with an average of $13.61$ golden chunks per query. Notably, the distribution of relevant documents is right-skewed, with a median of 6 and a high standard deviation ($\sigma=21.41$). This variance reflects the diverse nature of enterprise search, where some queries address specific technical identifiers with a single relevant source, while others address broad architectural topics with many relevant documentation fragments. To maintain the integrity of our public benchmark, we withhold the gold labels for the test queries to facilitate blind evaluation.
A.3 Design Rationale
Our pipeline reflects a principled balance between recall-oriented aggregation and precision-oriented filtering. The multi-retriever ensemble, comprising seven distinct models, ensures comprehensive coverage across the document corpus. By combining dense semantic encoders with lexical models (BM25), we capture complementary aspects of relevance that a single model might overlook. Following this "wide net" approach, LLM-based filtering leverages deep reasoning capabilities to adjudicate the final labels, eliminating false positives that exhibit surface-level similarity but fail to satisfy the semantic requirements of the query.
A.4 Analysis of Retriever Contributions
To validate the necessity of a multi-retriever ensemble, we investigate the individual contributions and potential redundancies of each model.
Individual Recall.
We first evaluate the standalone performance of each retriever. For a given query, we retrieve the top 420 candidates from a single model and calculate its recall against the final generated ground truth. As shown in Table 1, even the highest-performing model (gemini-embedding-001) achieves only 82.48% recall. This confirms that relying on a single retriever—no matter how powerful—would result in a significant loss of valid relevant documents, thereby biasing the dataset.
Redundancy and Diversity.
We further conduct a "leave-one-out" ablation study to determine if any retriever is redundant. For each model $M_{i}$, we aggregate 60 candidates from each of the other six retrievers (as we did during dataset generation) and measure the union’s recall against the ground truth. If the recall for a combination $\text{All}\setminus\{M_{i}\}$ were to reach 1.0, it would indicate that $M_{i}$ is redundant.As reported in Table 2, no model combination achieves perfect recall, with values ranging from 93.25% to 97.13%. This signifies that every retriever in our ensemble contributes unique relevant candidates that are not captured by the others. Notably, the drop in recall is most significant when removing the top-performing dense models, but even the removal of BM25 results in a loss of coverage, underscoring the necessity of a hybrid, multi-model approach for high-quality dataset construction.
Model Recall
gemini-embedding-001 82.48
gte-Qwen2-7B-instruct 82.25
SFR-Embedding-Mistral 79.20
text-embedding-3-large 75.54
Qwen3-Embedding-8B 70.12
embed-english-v3 65.83
BM25 52.18
Table 1: Recall@420 of Individual Retrievers on DevRev Search dataset
Model Combination Recall
All $\setminus$ {gemini-embedding-001} 93.25
All $\setminus$ {gte-Qwen2-7B-instruct} 95.86
All $\setminus$ {SFR-Embedding-Mistral} 96.30
All $\setminus$ {text-embedding-3-large} 97.13
All $\setminus$ {Qwen3-Embedding-8B} 96.83
All $\setminus$ {embed-english-v3} 95.61
All $\setminus$ {BM25} 95.96
Table 2: Leave-one-out ablation study on DevRev Search dataset
A.5 LLM-based Filtering Prompt
The prompt used for LLM-based filtering consists of annotation instructions, few-shot examples, and the target query-chunk pair. The full prompt is presented below.
A.5.1 System Prompt
Annotation Instructions
The focus is on whether an article chunk would help a support agent answer a query. Key instructions for annotators:
Focus on Problem in the query and Information in article chunk: Determine if the problem described in the query can be answered by the information present in article chunk. If article chunks information would likely answer the query, then article chunk should be labeled as relevant. For example, if the article chunk explains how to use certain features in the app and the query is also asking how to use those features (even if in different words).
IMPORTANT - Beware of Superficial Word Overlap: Do not label an article chunk as relevant only because it shares some keywords with the query. Read the article chunk and query fully - article chunk and query might both mention a common term (like "login") but could be about different aspects of login (one about UI for the login page, another about authentication). Only consider lexical overlap meaningful if the article chunk contains information to answer the query (e.g. the query asks how to solve a specific login issue, and the article chunk contains information to solve that specific login issue).
{few_shot_examples}
Edge case: In the case article chunk contains only partial information required to answer the query, label it relevant only in the case when it answers the query substantially. Simple lexical overlap does not imply relevance. When in doubt, ask: "Would a support agent benefit from seeing the article chunk while answering the query?" If yes, label it similar; if not, or only minimally, then its not relevant enough to help in the support workflow.
A.5.2 Few-Shot Examples
Examples of Relevant article chunk:
Example 1: Query: "Where can I find the DevRev API documentation?" and article chunk: "Resources to learn how to use DevRev APIs can be found at https://developer.devrev.ai/". The query is asking about where to find documentation on how to use DevRev APIs and the article chunk contains the information about the location where to find DevRev API documentation. The article chunk should be marked as relevant - it contains the information required to answer the query (even if words differ).
Example 2: Query: "What is a custom object?" and article chunk: "To create a custom object raise a support ticket. Custom objects are DevRev objects which can be customized". Even though the article chunk initially contains the information on how to create a custom object, it later also contains the information on what is a custom object which is what is asked in the query. Mark the article chunk relevant.
Examples of Non-Relevant article chunk:
Example 1: Query: "How to create a vista?" vs article chunk: "Vista is a list of DevRev objects" Both query and article chunk are about vistas and share the word "vista" but the information in article chunk is different from what query is asking about (query is asking how to create a vista, the article chunk is about what are vistas). This article chunk should be marked non relevant - information in the article chunk would not help answer the query.
Example 2: Query: "How to solve FORBIDDEN error when calling custom object API?" vs article chunk: "To solve BAD_REQUEST error when calling custom object API, look for DevRev custom object API documentation and fix your request structure". On the surface the article chunk looks relevant (same feature: custom object API). However, the error nature is different (one is a FORBIDDEN error, another is a BAD_REQUEST error). Unless further context in the article chunk reveals that both are the same errors, treat the article chunk as non relevant because the resolutions of both errors would differ (one might need more permissions, the other requires fixing the request).
[…]
A.5.3 User Prompt Template
Query: {query}
Article Chunk: {candidate}
A.6 Experimental Setup
We use AdamW optimizer with cosine learning rate scheduler.
Table 3 shows the hyperparameter configuration used by us.
Dataset Model Learning Rate Warmup Steps (%) LoRA Dropout (if LoRA)
DevRev Search snowflake-arctic-mbed-l-v2 $5\mathrm{e}{-6}$ 0 0.00
DevRev Search Qwen3-Embedding-4B $5\mathrm{e}{-6}$ 1010 0.05
Scifact snowflake-arctic-mbed-l-v2 $5\mathrm{e}{-6}$ 0 0.00
Scifact Qwen3-Embedding-4B $5\mathrm{e}{-6}$ 1010 0.05
Table 3: Hyperparameter settings for different datasets and embedding models