### v2

# Succeeding at Scale: Automated Dataset Construction and Query-Side Adaptation for Multi-Tenant Search

Abstract
Large-scale multi-tenant retrieval systems generate extensive query logs but lack curated relevance labels for effective domain adaptation, resulting in substantial underutilized “dark data”. This challenge is compounded by the high cost of model updates, as jointly fine-tuning query and document encoders requires full corpus re-indexing, which is impractical in multi-tenant settings with thousands of isolated indices. We introduce DevRev-Search [^2] https://huggingface.co/datasets/devrev/search, a passage retrieval benchmark for technical customer support built via a fully automated pipeline. Candidate generation uses fusion across diverse sparse and dense retrievers, followed by an LLM-as-a-Judge for consistency filtering and relevance labeling. We further propose an Index-Preserving Adaptation strategy that fine-tunes only the query encoder, achieving strong performance gains while keeping document indices fixed. Experiments on DevRev-Search, SciFact, and FiQA-2018 show that Parameter-Efficient Fine-Tuning (PEFT) of the query encoder delivers a remarkable quality–efficiency trade-off, enabling scalable and practical enterprise search adaptation.
1 Introduction
The transition from lexical matching methods such as BM25 (Robertson and Zaragoza (2009)) to dense neural retrieval has revolutionized information discovery (Karpukhin et al. (2020)). However, deploying bi-encoder architectures in multi-tenant enterprise environments presents a ”double scarcity” challenge. First, the Data Scarcity Bottleneck: enterprise tenants possess ”dark data,” proprietary corpora where relevance labels are non-existent, and standard benchmarks like BEIR (Thakur et al. (2021)) fail to capture the noisy, heterogeneous nature of these domains. Second, the Adaptation Latency Bottleneck: symmetric fine-tuning of both encoders (Qu et al. (2021)) incurs a massive ”Re-indexing Tax,” as any document encoder update necessitates re-generating embeddings for the entire corpus, computationally prohibitive for platforms hosting thousands of tenants.
To bridge these gaps, we present a unified methodology for scalable dataset construction and efficient model adaptation:
  1. 1.
    DevRev-Search Benchmark: A scalable pipeline that synthesizes training data without human annotators by pooling candidates from diverse retrievers and employing LLM-as-a-Judge filtering (Dai et al. (2023); Rahmani et al. (2024)).
  2. 2.
    Zero-Reindexing Adaptation: An asymmetric fine-tuning approach that adapts only the query encoder, enabling tenant-specific models with a frozen document index.
  3. 3.
    Parameter Efficient Query Adaptation: We show that parameter-efficient techniques such as LoRA (Hu et al. (2022)), embedding transformation using linear and feed-forward projections, and fine-tuning of only limited transformer layers—can achieve performance that matches or closely approaches full query adaptation. We further conduct ablations over LoRA ranks and modules, as well as the number of unfrozen transformer layers.
We validate our approach on DevRev-Search, SciFact (Wadden et al. (2022)), and FiQA-2018 (Maia et al. (2018)), demonstrating robust domain adaptation with a fraction of the cost of full bi-encoder fine-tuning.
2 Related Work
Synthetic Data: Addressing label scarcity, recent works leverage LLMs for synthetic data generation (Dai et al. (2023); Bonifacio et al. (2022); Wang et al. (2022)). A key component is consistency filtering, discarding queries that fail to retrieve their source. We extend this by employing fusion-based candidate generation (Cormack et al. (2009)), aggregating diverse retrievers to minimize single-model bias and improve coverage. LLM-as-Judge: While LLMs show promise as relevance assessors (Rahmani et al. (2024)), concerns regarding bias persist (Soboroff (2025)). We mitigate this by using LLMs primarily for filtering pooled candidates, identifying positives rather than generating them, and validating a subset with human annotators.
Index-Preserving Adaptation: Standard symmetric fine-tuning (Karpukhin et al. (2020)) incurs a prohibitive ”Re-indexing Tax”. Prior index-preserving approaches focus on pseudo-relevance feedback (Yu et al. (2021)) or asymmetric tuning (Wang and Lyu (2023)). We formalize this as Query-Only Adaptation, adapting the query manifold while freezing the document index. Parameter-Efficient Fine-Tuning (PEFT): PEFT methods like Adapters (Houlsby et al. (2019)), LoRA (Hu et al. (2022)), embedding projections (Yoon et al. (2024)) and partial unfreezing of layers (Lee et al. (2022)) have proven effective in retrieval (Litschko et al. (2022)). Unlike full fine-tuning, PEFT methods use much fewer trainable parameters compared to full fine-tuning. Building on findings regarding intrinsic dimensionality (Aghajanyan et al. (2020)), we demonstrate that applying PEFT to the query encoder results in both high quality and efficiency for multi-tenant serving.
3 Dataset Generation
Recent retrieval gains are driven by high-quality, domain-specific data, yet publicly available enterprise search datasets remain scarce. Existing benchmarks such as MS-MARCO and SciFact focus on web passages or scientific abstracts and do not capture the semi-structured, heterogeneous nature of enterprise data, including support tickets, issue trackers, and internal documentation. DevRev-Search addresses this gap by providing a high-fidelity benchmark for enterprise-specific retrieval.
Manual annotation is costly, slow, and inherently low-recall, as annotators cannot exhaustively review large corpora, resulting in systematic false negatives. To address this, we introduce a scalable, automated dataset construction pipeline that maximizes candidate coverage through multi-stage retrieval while maintaining high precision via an LLM-as-judge.
3.1 Query Collection and Cleaning
We collected customer queries from production agent interactions as our source of real-world question data. However, raw customer queries often contain noise, including test queries, code snippets, and malformed inputs that are not legitimate natural language questions. To ensure dataset quality, we implemented a multi-stage filtering process: (1) Length filtering: Removing queries with word counts in the bottom and top 25% percentiles. (2) Language detection: Retaining only English queries. (3) Deduplication: Removing exact duplicate queries. (4) Clustering-based diversity: Selecting representative samples from clusters to ensure diversity and avoid semantic repetition.
3.2 Document Segmentation and Semantic Granularity
Enterprise documents are typically long and sparsely relevant to specific queries, which poses challenges for dense retrieval due to encoder token limits and the limited expressive capacity of fixed-size embeddings. Encoding entire documents often results in diluted representations that obscure fine-grained relevance. To address this, we collected all public documentation of DevRev and applied Recursive Character Splitting LangChain (2025). Documents are segmented into chunks of up to 500 characters with no overlap, maximizing semantic coverage while preserving precision. The recursive strategy favors natural structural boundaries (e.g., paragraphs, sentences, then spaces), ensuring each fragment is both embedding-friendly and semantically self-contained.
3.3 Multi-Retriever Annotation
To create high-quality query-document pairs, we employed an ensemble retrieval approach designed to maximize recall while maintaining precision.
Retrieval Ensemble: We applied an ensemble of seven diverse models: six dense retrievers (gemini-embedding-001 Google (2025), text-embedding-3-large (OpenAI (2024)), embed-english-v3 (Cohere (2023)), Qwen-3-Embedding-8B (Zhang et al. (2025)), GTE-Qwen2-7B-Instruct (Li et al. (2023)), SFR-Embedding-Mistral (Meng et al. (2024))) and one lexical retriever (BM25) ( Robertson and Zaragoza (2009)), each returning the top 60 document chunks.
Union-based Aggregation: We computed the union of results from all retrievers to create a comprehensive candidate set of potentially relevant chunks. This ensures that chunks retrieved by any of the seven models are included in the candidate set, maximizing coverage and recall across different retrieval paradigms yielding $\geq 60$ and $\leq 420$ unique candidate chunks per query across all models.
LLM-based Filtering: While the ensemble approach boosts recall, it introduces noise via false positives. To improve precision, we applied LLM-based filtering to the fused candidate set. Using a carefully designed prompt (Appendix A.3), a large language model was tasked with retaining only document chunks genuinely relevant to each query, removing those with high lexical or semantic similarity but lacking substantive answer content.
Quality Validation: To verify the reliability of our automated annotation process, we randomly sampled $10\%$ queries and manually validated the final annotations. This validation confirmed the accuracy of our annotation pipeline.
3.3.1 Analysis of Retriever Contributions
To assess the necessity of a multi-retriever ensemble, we analyzed the individual performance, overlap, and complementarity of each retriever.
Individual Recall.
We measured standalone recall by retrieving the top 420 candidates per query from each model and comparing them to the final ground truth. As shown in Table 3 (Appendix A.2), even the best-performing retriever (gemini-embedding-001) achieves only 82.48 recall, showing that any single retriever misses a substantial portion of relevant documents and introduces bias.
Redundancy and Diversity.
To evaluate redundancy, we performed a leave-one-out ablation: for each retriever $M_{i}$, we aggregated candidates from the remaining six models and measured recall against the ground truth. If $\text{All}\setminus{M_{i}}$ reached perfect recall, $M_{i}$ would be redundant. However, as reported in Table 4 (Appendix A.2), recall ranges from 93.25 to 97.13, demonstrating that every retriever contributes unique relevant candidates. While removing strong dense models causes the largest drop, excluding BM25 also reduces coverage, highlighting the value of a hybrid ensemble.
Overall, our pipeline balances recall-oriented aggregation with precision-oriented filtering. The seven-retriever ensemble provides broad coverage by combining dense semantic and lexical signals, while LLM-based filtering applies deep reasoning to eliminate false positives and finalize high-quality labels. See Appendix A.1 for additional dataset statistics.
4 Query-Only Adaptation
Standard bi-encoder fine-tuning updates both the query ($E_{q}$) and document ($E_{d}$) encoders, but modifying $E_{d}$ requires costly re-embedding and index reconstruction (e.g., HNSW) over millions of documents, which is impractical in large-scale, multi-tenant systems. To eliminate this re-indexing cost while enabling rapid iteration, we propose Query-Only Adaptation, which freezes the document encoder and index for very fast deployment.
We evaluate this approach across three datasets with varying relevance densities: DevRev-Search (enterprise, high density; 13.6 relevant chunks/query), SciFact (scientific, low density; 1.1), and FiQA-2018 (financial, medium density; 2.6). These respectively emphasize recall, precise evidence retrieval, and balanced precision–recall. We use snowflake-arctic-embed-l-v2.0 Yu et al. (2024) and Qwen3-Embedding-4B Zhang et al. (2025) as backbone encoders.
Training employs InfoNCE loss with mined hard negatives to improve discrimination among semantically similar documents. A single tuned temperature is used throughout training, and stability is ensured via asynchronous ANCE training (Xiong et al., 2020) to prevent representation collapse. Additional experimental details are provided in Appendix A.4.
Our analysis focuses on: (1) Query-Only vs. Query–Document Fine-Tuning, measuring the performance impact of freezing the document index; and (2) Parameter-Efficient Query Adaptation, evaluating whether low-rank and parameter-efficient methods can match full fine-tuning while enabling cost-effective multi-tenant deployment.
4.1 Query Only vs Query-Document Fine-Tuning
We first evaluate whether asymmetric Query-Only fine-tuning ($Q$) can reach performance parity with symmetric Query-Document fine-tuning ($QD$). Table 1 summarizes the results across both architectures. The data demonstrates that freezing the document encoder incurs minimal performance loss. In all cases, the $Q$ strategy significantly improves over the untrained baseline. Crucially, it remains highly competitive with the computationally expensive $QD$ approach, achieving comparable retrieval metrics across all domains and, in the case of qwen3-4b on SciFact, even marginally outperforming the joint tuning strategy in Recall@10.
Table 1: Comparison of Recall@10 and NDCG@10 in retrieval performance for baseline untrained model (Base) with Query-Document (QD), and Query-Only (Q) finetuning
Model Variant DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Base 0.256 0.304 0.819 0.691 0.533 0.459
QD 0.314 0.362 0.869 0.748 0.578 0.505
Q 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Base 0.219 0.264 0.903 0.769 0.649 0.571
QD 0.325 0.396 0.949 0.875 0.712 0.628
Q 0.327 0.370 0.953 0.830 0.688 0.602
4.2 Parameter Efficient Fine-Tuning
We examine whether Query-Only full fine-tuning performance can be achieved using PEFT methods, reducing trainable parameters and enabling scalable multi-tenant adaptation. We evaluate several PEFT techniques: LoRA, a linear projection head, a 2 hidden-layer feed-forward network (FFN) on embeddings, and unfreezing the top 8 transformer layers. The number and proportion of trainable parameters for each method are reported in Table 5 (Appendix A.4).
Results in Table 2 show that the best LoRA configuration consistently matches or outperforms full fine-tuning, likely due to its implicit regularization via layer-wise residual learning. Linear and FFN heads are also competitive; notably, FFN surpasses full fine-tuning on SciFact. In contrast, fine-tuning only the top 8 transformer layers underperforms despite using more parameters, highlighting LoRA’s effectiveness as a PEFT strategy.
Table 2: Comparison of retrieval performance for various parameter-efficient methods with full finetuning for query-only adaptation (Here *Lora and 8-Tr refer to the best LoRA configuration and fine-tuning only 8 transformer layers at the top of the base model, respectively)
Model Type DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Linear 0.271 0.314 0.849 0.722 0.543 0.470
FFN 0.281 0.318 0.880 0.743 0.534 0.464
*LoRA 0.309 0.342 0.915 0.827 0.555 0.478
8-Tr 0.273 0.317 0.849 0.718 0.547 0.476
Full 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Linear 0.331 0.358 0.935 0.815 0.666 0.583
FFN 0.340 0.362 0.946 0.861 0.660 0.577
*LoRA 0.355 0.399 0.950 0.844 0.694 0.607
8-Tr 0.312 0.356 0.932 0.802 0.673 0.590
Full 0.327 0.370 0.953 0.830 0.688 0.602
Figure 1: Ablation results for query only adapation for DevRev-Search across (a) varying LoRA ranks (with $\alpha=r$), (b) different combinations of target LoRA modules, and (c) varying number of top unfrozen transformer layers with the rest of the model frozen
LoRA Rank Ablations.
We study sensitivity to the LoRA rank ($r$) with $\alpha=r$. On DevRev-Search (Fig. 1(a)), arctic-l-v2 improves monotonically and saturates at high ranks, whereas qwen3-4b shows a peaked profile with best performance at $r=32$, likely due to overfitting at larger ranks. Results on SciFact and FiQA (Table 7, Appendix A.5) show a consistent pattern: $r\in[32,64]$ provides the best trade-off between capacity and regularization. Although $r=128$ can sometimes yield minor gains, performance largely saturates, making $32$ or $64$ the most efficient choices.
Lora Target Module Ablations.
We also compare LoRA-based query-only adaptation across transformer sub-layer combinations — dense, QV (Query-Value), QKV (Query-Key-Value), and all modules. Figure 1(b) shows the DevRev-Search results. For results on the other datasets, please refer to Table 8 (Appendix A.5). While generally taking all layers is the best choice, we observe that just tuning dense layers provides an impressive performance-efficiency tradeoff compared to other modules. This observation is more visible with qwen3-4b.
Transformer Layer Ablations.
Finally, we freeze the base model and ablate the number of transformer layers at the top that are unfrozen. As seen in Fig 1(c), the performance monotonically increases with the number of unfrozen transformer layers, regardless of model architecture, on DevRev-Search. Table 9 (Appendix A.5) reports the results for the other datasets, which exhibit the same trend.
5 Conclusion
We presented a unified approach to scalable dataset construction and efficient model adaptation for multi-tenant retrieval. Our automated pipeline, combining multi-retriever fusion with LLM-based consistency filtering, produces high-quality training data without manual annotation, yielding the DevRev-Search benchmark. Our index-preserving adaptation strategy, fine-tuning only the query encoder via PEFT, achieves competitive performance while eliminating re-indexing overhead, thus optimizing for the quality-efficiency trade-off.
Appendix A Appendix
A.1 Dataset Statistics
We partition the DevRev-Search dataset into a training set and an evaluation (test) set. The dataset consists of 291 training queries and 92 test queries. The training set exhibits a rich density of relevant documents, with an average of $13.61$ golden chunks per query. Notably, the distribution of relevant documents is right-skewed, with a median of 6 and a high standard deviation ($\sigma=21.41$). This variance reflects the diverse nature of enterprise search, where some queries address specific technical identifiers with a single relevant source, while others address broad architectural topics with many relevant documentation fragments. To maintain the integrity of our public benchmark, we withhold the gold labels for the test queries to facilitate blind evaluation.
A.2 Retriever Contributions
Table 3: Recall@420 of Individual Retrievers on DevRev-Search dataset
Model Recall
gemini-embedding-001 82.48
gte-Qwen2-7B-instruct 82.25
SFR-Embedding-Mistral 79.20
text-embedding-3-large 75.54
Qwen3-Embedding-8B 70.12
embed-english-v3 65.83
BM25 52.18
Table 4: Leave-one-out ablation study on DevRev-Search dataset
Model Combination Recall
All $\setminus$ {gemini-embedding-001} 93.25
All $\setminus$ {gte-Qwen2-7B-instruct} 95.86
All $\setminus$ {SFR-Embedding-Mistral} 96.30
All $\setminus$ {text-embedding-3-large} 97.13
All $\setminus$ {Qwen3-Embedding-8B} 96.83
All $\setminus$ {embed-english-v3} 95.61
All $\setminus$ {BM25} 95.96
A.3 LLM-based Filtering Prompt
The prompt used for LLM-based filtering consists of annotation instructions, few-shot examples, and the target query-chunk pair. The full prompt is presented below.
A.3.1 System Prompt
Annotation Instructions
The focus is on whether an article chunk would help a support agent answer a query. Key instructions for annotators:
Focus on Problem in the query and Information in article chunk: Determine if the problem described in the query can be answered by the information present in article chunk. If article chunks information would likely answer the query, then article chunk should be labeled as relevant. For example, if the article chunk explains how to use certain features in the app and the query is also asking how to use those features (even if in different words).
IMPORTANT - Beware of Superficial Word Overlap: Do not label an article chunk as relevant only because it shares some keywords with the query. Read the article chunk and query fully - article chunk and query might both mention a common term (like "login") but could be about different aspects of login (one about UI for the login page, another about authentication). Only consider lexical overlap meaningful if the article chunk contains information to answer the query (e.g. the query asks how to solve a specific login issue, and the article chunk contains information to solve that specific login issue).
{few_shot_examples}
Edge case: In the case article chunk contains only partial information required to answer the query, label it relevant only in the case when it answers the query substantially. Simple lexical overlap does not imply relevance. When in doubt, ask: "Would a support agent benefit from seeing the article chunk while answering the query?" If yes, label it similar; if not, or only minimally, then its not relevant enough to help in the support workflow.
A.3.2 Few-Shot Examples
Examples of Relevant article chunk:
Example 1: Query: "Where can I find the DevRev API documentation?" and article chunk: "Resources to learn how to use DevRev APIs can be found at https://developer.devrev.ai/". The query is asking about where to find documentation on how to use DevRev APIs and the article chunk contains the information about the location where to find DevRev API documentation. The article chunk should be marked as relevant - it contains the information required to answer the query (even if words differ).
Example 2: Query: "What is a custom object?" and article chunk: "To create a custom object raise a support ticket. Custom objects are DevRev objects which can be customized". Even though the article chunk initially contains the information on how to create a custom object, it later also contains the information on what is a custom object which is what is asked in the query. Mark the article chunk relevant.
Examples of Non-Relevant article chunk:
Example 1: Query: "How to create a vista?" vs article chunk: "Vista is a list of DevRev objects" Both query and article chunk are about vistas and share the word "vista" but the information in article chunk is different from what query is asking about (query is asking how to create a vista, the article chunk is about what are vistas). This article chunk should be marked non relevant - information in the article chunk would not help answer the query.
Example 2: Query: "How to solve FORBIDDEN error when calling custom object API?" vs article chunk: "To solve BAD_REQUEST error when calling custom object API, look for DevRev custom object API documentation and fix your request structure". On the surface the article chunk looks relevant (same feature: custom object API). However, the error nature is different (one is a FORBIDDEN error, another is a BAD_REQUEST error). Unless further context in the article chunk reveals that both are the same errors, treat the article chunk as non relevant because the resolutions of both errors would differ (one might need more permissions, the other requires fixing the request).
[…]
A.3.3 User Prompt Template
Query: {query}
Article Chunk: {candidate}
A.4 Experimental Setup
All our trainings optimize InfoNCE loss using AdamW optimizer, with 16 mined hard negatives per query, with 8 negatives randomly sampled from them each time. A learning rate of $5\times 10^{-6}$ is used for all runs with a cosine scheduler and no warm-up. For all the experiments, we use a uniform temperature of 0.1, which was found to work the best.
Additionally, we noticed that simply using fixed mined hard negatives leads to cases of representation collapse wherein InfoNCE loss keeps improving, but training metrics rapidly deteriorate after a point, especially in datasets like FiQA and DevRev-Search. Upon further analysis, we noticed that this happens due to evolving hard negatives, where over the course of training, old hard negatives become ’simple’ and new hard negatives replace them. To counteract this, we adopt the asynchronous ANCE training proposed by Xiong et al. (2020) to periodically update the hard negatives in InfoNCE loss at every 200 steps. This proved to improve the performance in most cases, in addition to much more stable training as shown in Fig 2. Quantitative comparison of results obtained with and without ANCE is shown in Table 6
The feed-forward network (FFN) used in our PEFT experiments consists of GELU (Hendrycks (2016)) activations and 2 hidden layers, each with the same number of nodes as the base model’s embedding dimension. Also, for both linear and FFN heads, we find that zeroing biases and initializing weight matrices to the identity perform consistently better than standard initializations such as Xavier (Glorot and Bengio (2010)) or Kaiming (He et al. (2015)).
Table 5: Parameter efficiency comparison for Query-Only adaptation across LoRA ranks, unfrozen layers, and head-only training.
Model Configuration Trainable Total % Full
Params Params
arctic-l-v2 Full Model 566,705,152 566,705,152 100.00
LoRA $r=8$ 3,538,944 570,244,096 0.62
LoRA $r=16$ 7,077,888 573,783,040 1.25
LoRA $r=32$ 14,155,776 580,860,928 2.50
LoRA $r=64$ 28,311,552 595,016,704 4.99
LoRA $r=128$ 56,623,104 623,328,256 9.99
Linear Head (Frozen Base) 1,049,600 567,754,752 0.19
MLP Head (Frozen Base) 3,148,800 569,853,952 0.56
1 Unfrozen Layer 12,596,224 566,705,152 2.22
2 Unfrozen Layers 25,192,448 566,705,152 4.45
4 Unfrozen Layers 50,384,896 566,705,152 8.89
8 Unfrozen Layers 100,769,792 566,705,152 17.78
qwen3-4b Full Model 4,021,774,336 4,021,774,336 100.00
LoRA $r=8$ 16,515,072 4,038,289,408 0.41
LoRA $r=16$ 33,030,144 4,054,804,480 0.82
LoRA $r=32$ 66,060,288 4,087,834,624 1.64
LoRA $r=64$ 132,120,576 4,153,894,912 3.29
LoRA $r=128$ 264,241,152 4,286,015,488 6.57
Linear Head (Frozen Base) 6,556,160 4,028,330,496 0.16
MLP Head (Frozen Base) 19,668,480 4,041,442,816 0.49
1 Unfrozen Layer 100,933,376 4,021,774,336 2.51
2 Unfrozen Layers 201,864,192 4,021,774,336 5.02
4 Unfrozen Layers 403,725,824 4,021,774,336 10.04
8 Unfrozen Layers 807,449,088 4,021,774,336 20.08
Table 6: Retrieval performance with and without ANCE for full Query-Only adaptation
Model DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Standard 0.291 0.331 0.860 0.731 0.543 0.473
+ANCE 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Standard 0.333 0.373 0.941 0.813 0.679 0.593
+ANCE 0.327 0.370 0.953 0.830 0.688 0.602
Figure 2: Training NDCG@10 vs training step for Query-Only adaptation of arctic-l-v2 (full model)
A.5 Ablation Results
Table 7: Retrieval performance variation across different LoRA ranks ($r=\alpha$ for all configurations)
Model LoRA Rank DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 8 0.258 0.307 0.867 0.738 0.540 0.469
16 0.268 0.322 0.880 0.748 0.548 0.474
32 0.283 0.328 0.905 0.769 0.554 0.478
64 0.304 0.341 0.914 0.807 0.555 0.478
128 0.309 0.342 0.915 0.827 0.555 0.478
qwen3-4b 8 0.304 0.346 0.948 0.826 0.685 0.603
16 0.329 0.360 0.948 0.832 0.686 0.605
32 0.355 0.399 0.950 0.844 0.690 0.607
64 0.336 0.369 0.950 0.816 0.694 0.607
128 0.336 0.380 0.957 0.831 0.682 0.601
Table 8: Retrieval performance variation across different target modules for LoRA $r=64$ and $\alpha=64$
Model Layers DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Dense 0.287 0.329 0.854 0.726 0.551 0.477
QV 0.283 0.325 0.854 0.726 0.554 0.476
QKV 0.283 0.325 0.854 0.726 0.554 0.477
All 0.304 0.341 0.914 0.807 0.555 0.478
qwen3-4b Dense 0.333 0.371 0.950 0.818 0.693 0.607
QV 0.324 0.377 0.923 0.793 0.688 0.605
QKV 0.337 0.368 0.930 0.796 0.689 0.606
All 0.336 0.369 0.950 0.816 0.694 0.607
Table 9: Retrieval performance variation across different numbers of unfrozen transformer layers at the top of the base model, with the rest of the layers frozen
Model Layers DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 1 0.258 0.306 0.827 0.700 0.534 0.462
2 0.260 0.306 0.835 0.709 0.536 0.464
4 0.267 0.311 0.835 0.712 0.539 0.471
8 0.273 0.317 0.849 0.718 0.547 0.476
qwen3-4b 1 0.237 0.290 0.903 0.778 0.661 0.576
2 0.259 0.306 0.908 0.783 0.664 0.579
4 0.275 0.325 0.912 0.788 0.667 0.583
8 0.312 0.356 0.932 0.802 0.673 0.590